OpenStudy (anonymous):

determine the intercept 9x-6y+3=0

OpenStudy (kc_kennylau):

Put y=0 to obtain the x-intercept and x=0 the y-intercept.

OpenStudy (anonymous):

9x-6y+3=0 9*0+6y+3=0 0+6y+3=0 6y+3-3=0-3 6y/6=-3/6 y=-1/2 is this accurate? (0,-1/2)

OpenStudy (kc_kennylau):

Don't turn the -6y into +6y in the second step !!!

OpenStudy (anonymous):

it should be -6y+3-3?

OpenStudy (kc_kennylau):

Therefore the answer is?